The sum of the 4th and 6th term of an arithmetical progression is 42.the sum of the 3rd and 9th terms of the progression is 52.find the first term,the common difference and the sum of the first ten terms of the progression.
Answer by mananth(16946)

The sum of the 4th and 6th term of an arithmetical progression is 42.the sum of the 3rd and 9th terms of the progression is 52.find the first term,the common difference and the sum of the first ten terms of the progression.
T4+t6=42
a+3d+a+5d=42
2a+8d=42
a+4d=21....................(1)
t3+t9=52
a+2d+a+8d=52
2a+10d=52
a+5d=26....................(2)
subtract (2) from (1)
-d=-5
d=5
a+4d=21
a+20=21
a=1
a=1,d=5
1,6,11,16,21,26
t4+t6=42
You find S10